OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 15x – y = –6 5x – 3y = –13 5(–15x – 6) – 3y = –13 5(15x + 6) – 3y = –13 5x – 3(15x + 6) = –13 5x – 3(–15x – 6) = –13

OpenStudy (anonymous):

5x -3(15x+6)=-13

OpenStudy (anonymous):

15x-y=-6. -y=-15x-6. y=15x+6.

OpenStudy (anonymous):

thankyou so much.... XD